### v2

# Adaptive Query Algorithms for Relational Structures Based on Homomorphism Counts

Abstract
A query algorithm based on homomorphism counts is a procedure to decide membership for a class of finite relational structures using only homomorphism count queries. A left query algorithm can ask the number of homomorphisms from any structure to the input structure and a right query algorithm can ask the number of homomorphisms from the input structure to any other structure. We systematically compare the expressive power of different types of left or right query algorithms, including non-adaptive query algorithms, adaptive query algorithms that can ask a bounded number of queries, and adaptive query algorithms that can ask an unbounded number of queries. We also consider query algorithms where the homomorphism counting is done over the Boolean semiring $\mathbb{B}$, meaning that only the existence of a homomorphism is recorded, not the precise number of them.
keywords:
Query algorithms, homomorphisms, homomorphism counts, directed graphs, relational structures, Datalog, constraint satisfaction
1 Introduction
A classic result by Lovász [lovasz1967operations] states that a finite relational structure $A$ is determined up to isomorphism by the numbers $\hom(B,A)$ of homomorphisms from $B$ to $A$ for every finite relational structure $B$ with the same signature. This list of numbers is called the left homomorphism profile of $A$ and is written as $\hom(\mathcal{C},A)$, where $\mathcal{C}$ is the class of all finite relational structures with the same signature as $A$. Similarly, one can define the right homomorphism profile as the list of numbers of homomorphism from $A$ to each structure. The right profile also determines the structure up to isomorphism, as shown by Chaudhuri and Vardi [chaudhuri1993optimization].
Recently, there has been considerable interest in exploring which structures can be distinguished by the homomorphism profiles when the class $\mathcal{C}$ is restricted. As an example, in 2010, Dvořák [dvovrak2010recognizing] proved that if $\mathcal{C}$ is the class of graphs of treewidth at most $k$, the left profile can distinguish precisely the same graphs as the $k$-dimensional Weisfeiler-Leman algorithm, a well-known polynomial time graph isomorphism test. Cai, Fürer, and Immerman [cai1992optimal] had already proved that the $k+1$ variable fragment of first-order logic extended with counting quantifiers can distinguish precisely the same graphs as $k$-dimensional Weisfeiler-Leman algorithm. Also, in 2020, Grohe [grohe2020counting] proved that first-order logic with counting of quantifier rank at most $k$ can distinguish exactly the same graphs as the left homomorphism profile restricted to graphs of treedepth $k$. Mančinska and Roberson [Mancinska2020:quantum] showed that the graphs that are quantum isomorphic are exactly the graphs that have the same left homomorphism profile restricted to planar graphs. In 2024, Seppelt [Seppelt2024:logical] provides further insight into the logical equivalence relations for graphs that can be characterized by restricted left homomorphism profiles, using techniques from structural graph theory. These results highlight the importance of homomorphism profiles in finite model theory by describing a close connection between homomorphism profiles and logical languages.
Taking a slightly different approach, in 2022, Chen et al. [chen_et_al:LIPIcs.MFCS.2022.32] ask what properties of graphs can be decided based on finitely many homomorphism count queries. They did this by introducing the notion of a homomorphism-count based query algorithm. A non-adaptive left $k$-query algorithm (also referred to as a left $k$-query algorithm) consists of a finite tuple $\mathcal{F}=(F_{1},\ldots,F_{k})$ of structures and a subset $X\subseteq\mathbb{N}^{k}$. The algorithm is said to decide the class of structures
$\mathcal{A}\coloneqq\{A:(\hom(F_{i},A))_{i=1}^{k}\in X\}.$
Non-adaptive right $k$-query algorithms are defined similarly, but using $\hom(A,F_{i})$ instead of $\hom(F_{i},A)$. These notions allow for the study of what classes of structures can be defined using a finite restriction of the left (respectively, right) homomorphism profile. Chen et al. also define an adaptive version of left/right $k$-query algorithms. These are, roughly speaking, algorithms that can ask $k$ homomorphism-count queries where the choice of the $(i+1)$-th query may depend on the answers to the first $i$ queries. Chen et al. [chen_et_al:LIPIcs.MFCS.2022.32] studied the expressive power of non-adaptive and adaptive query algorithms for the specific case of simple graphs (i.e., undirected graphs without self-loops). Among other things, they showed that every first-order sentence $\phi$ that is a Boolean combination of universal first-order sentences can be decided by a non-adaptive $k$-left query algorithm for some $k$. On the other hand, they showed that there are first-order sentences $\phi$ that cannot be decided by a non-adaptive left $k$-query algorithm for any $k$. When it comes to adaptive query algorithms, they showed that every class of simple graphs can be decided by an adaptive left 3-query algorithm, but that there exists a class of simple graphs that is not decided by an adaptive right $k$-query algorithm for any $k$, when using only simple graphs for the queries.
In a 2024 paper, ten Cate et al. [ten2024homomorphism] extend the above analysis by exploring query algorithms for arbitrary relational structures (not only simple graphs) and by considering also query algorithms over the Boolean semiring $\mathbb{B}$, which means that instead of being able to query for the number of homomorphisms, the algorithm can only query for the existence of a homomorphism. The query algorithms as defined by Chen et al. [chen_et_al:LIPIcs.MFCS.2022.32] can be viewed as query algorithms over the semiring $\mathbb{N}$ of natural numbers. It is shown in [ten2024homomorphism] that, for classes of structures that are closed under homomorphic equivalence, non-adaptive left $k$-query algorithms (over $\mathbb{N}$) are no more powerful than non-adaptive left $k$-query algorithms over $\mathbb{B}$. In other words, “counting does not help” for such classes. Moreover, for such classes, non-adaptive left query algorithms are equivalent in expressive power to first-order logic, meaning that such classes are first-order definable if and only if they are decided by a non-adaptive left query algorithm over $\mathbb{N}$ (or $\mathbb{B}$).
[…]
Figure 1: Summary of our results regarding the relative expressive power of different types of left query algorithms over $\mathbb{N}$. Here, $\mathfrak{N}_{k}$, $\mathfrak{A}k$, and $\mathfrak{A}{unb}$ denote the collections of classes of structures that admit a non-adaptive left $k$-query algorithm, an adaptive left $k$-query algorithm, and an adaptive left unbounded query algorithm, respectively. No inclusions hold other than those depicted. In particular, $\mathfrak{A}_{2}\not\subseteq\bigcup_{k}\mathfrak{N}_{k}$ and $\mathfrak{N}_{k+1}\nsubseteq\mathfrak{A}_{k}$.
Our contributions.
We study the expressive power of different versions of these query algorithms for classes of relational structures. In doing so, we characterize their expressive power, relate it to that of logics and database query languages, and compare it across different types of algorithms. We focus mainly on adaptive query algorithms, including the previously unexplored notion of adaptive unbounded query algorithms—i.e., adaptive query algorithms for which there is no uniform bound on the number of queries they can ask on a given input, but which nevertheless terminate on every input.
Our main contributions are along two lines:
  1. 1.
    We exploit the symmetry of directed cycles and cycle-like structures to give a simple formula (Lemma 4.2) for the number of homomorphisms from a structure to a disjoint union of cycle-like structures of the same size. This formula serves as a powerful tool for deriving homomorphism indistinguishability results for finite profiles, and may be useful more broadly in related contexts. We use it to systematically investigate and compare the expressive power of adaptive and non-adaptive query algorithms over $\mathbb{N}$ for relational structures. For instance, we show, in Theorem LABEL:better-corollary, that for every signature containing a non-unary relation, there exists a class of structures of that signature that is not decided by an adaptive left $k$-query algorithm over $\mathbb{N}$ for any $k$. A concrete example of such a class is the class of directed graphs whose number of connected components is an even power of 2. This result is in sharp contrast to the result of Chen et al., which states that every class of simple graphs can be decided by an adaptive left 3-query algorithm over $\mathbb{N}$. We also show in Theorem LABEL:adaptive-not-better that, for each $k>1$, there exists a class that is decided by a non-adaptive left $k$-query algorithm over $\mathbb{N}$ but not by any adaptive left $(k-1)$-query algorithm over $\mathbb{N}$. However, in Theorem LABEL:A2_not_non-adaptive we also note that there exists a class that is decided by an adaptive left 2-query algorithm but not by any non-adaptive left query algorithm. We thus establish all inclusion and non-inclusion relations between the expressive powers of adaptive and non-adaptive left $k$-query algorithms over $\mathbb{N}$ (note: it follows from the proof of Lovász’s theorem that every class is decided by an adaptive left unbounded query algorithm). The results are summarized in Figure 1.
    Interestingly, the situation is quite different for right query algorithms over $\mathbb{N}$. Indeed, it follows from results in [wu2023study] that every class of structures can be decided with only two adaptive right queries. This is again in contrast to a prior result of Chen et al. [chen_et_al:LIPIcs.MFCS.2022.32], namely that there is no $k$ such that every class of simple graphs can be decided by an adaptive right $k$-query algorithm (when the queries themselves are also required to be simple graphs).
  2. 2.
    We investigate the expressive power of adaptive query algorithms over $\mathbb{B}$. We observe that adaptive bounded query algorithms over $\mathbb{B}$ have the same expressive power as non-adaptive query algorithms over $\mathbb{B}$, whose expressive power was characterized by ten Cate et al. [ten2024homomorphism]. However, we show that adaptive unbounded query algorithms over $\mathbb{B}$ are strictly more expressive. The classes decided by such algorithms exhibit an intricate pattern, which we show can be captured by a topological characterization. In the case of homomorphism-closed classes, we further show that the characterization can be simplified significantly, yielding an elegant condition for establishing whether such a class is decided by an adaptive unbounded query algorithm over $\mathbb{B}$ or not.
    We also apply the topological characterization to examine the expressive power in the special cases of classes that are homomorphism equivalence types, CSPs, or Datalog-definable. Specifically, we show that unboundedness does not increase the expressive power of adaptive left query algorithms over $\mathbb{B}$ for deciding CSPs and homomorphic-equivalence types, thereby yielding, based on existing results, an effective criterion for determining whether such classes are decided by an adaptive left unbounded query algorithm over $\mathbb{B}$. Additionally, we characterize the Datalog definable classes that are decided by such an algorithm in terms of their upper envelopes. Finally, we show that analogous results hold for adaptive right unbounded query algorithms over $\mathbb{B}$.
Outline.
Section 2 reviews preliminary definitions and background, including those of relational structures, digraphs, and walks in digraphs. Section 3 introduces the main concepts studied in this paper. We review the definition of non-adaptive query algorithms and introduce a new definition of adaptive query algorithms that applies in both the bounded and unbounded settings. Section 4 contains our main results regarding the expressive power of query algorithms over $\mathbb{N}$. Section 4.1 begins with the statement of a key lemma and then builds towards the result that not all classes can be expressed by an adaptive left bounded query algorithm over $\mathbb{N}$. In Section LABEL:sec:adaptive_vs_non-adaptive, we establish the results comparing adaptive and non-adaptive algorithms (as depicted in Figure 1). Section LABEL:right-algs-over-N concisely describes the simpler situation for adaptive right query algorithms over $\mathbb{N}$. Section LABEL:algs-over-B presents our results regarding the expressive power of query algorithms over $\mathbb{B}$. In Section LABEL:sec:unb_helps_B, we show that unboundedness increases the expressive power of adaptive left query algorithms over $\mathbb{B}$. Section LABEL:Unbounded_qals_over_B provides the topological characterization of the classes expressed by adaptive left unbounded query algorithms over $\mathbb{B}$, and in Section LABEL:sec:case_studies we apply this characterization to special cases. In Section LABEL:sec:rqa_over_B, we show that similar results can be obtained for adaptive right query algorithms over $\mathbb{B}$. We conclude in Section LABEL:sec:discussion by listing directions for further research. In the interest of space, we have moved the proofs of some theorems to the appendices.
2 Preliminaries
Relational structures and homomorphism counts.
A signature $\tau=\{R_{1},\ldots,R_{n}\}$ is a set of relational symbols where each symbol $R_{i}$ has an associated arity $r_{i}$. A relational structure $\mathbf{A}$ of signature $\tau$ consists of a non-empty set $A$ and a relation $R_{i}^{\mathbf{A}}\subseteq A^{r_{i}}$ for each symbol $R_{i}\in\tau$. We use $\mathsf{FIN}(\tau)$ to denote the class of all finite relational structures with signature $\tau$. We will only study finite relational structures; we will thus write “structure” and mean “finite relational structure” throughout this text. If $\mathbf{A}=(A,(R_{i}^{\mathbf{A}})_{i\in\{1,\ldots,n\}})$ and $\mathbf{B}=(B,(R_{i}^{\mathbf{B}})_{i\in\{1,\ldots,n\}})$ are structures with signature $\tau$, we say that $\varphi:A\to B$ is a homomorphism from $\mathbf{A}$ to $\mathbf{B}$ if it preserves all the relations, i.e. if
$(a_{1},\ldots,a_{r_{i}})\in R_{i}^{\mathbf{A}}\Longrightarrow(\varphi(a_{1}),\ldots,\varphi(a_{r_{i}}))\in R_{i}^{\mathbf{B}}$
for each $R_{i}\in\tau$. We then write $\varphi:\mathbf{A}\to\mathbf{B}$. If there exists a homomorphism from $\mathbf{A}$ to $\mathbf{B}$ we write $\mathbf{A}\to\mathbf{B}$, otherwise, $\mathbf{A}\nrightarrow\mathbf{B}$. If $\mathbf{A}\to\mathbf{B}$ and $\mathbf{B}\to\mathbf{A}$, we then write $\mathbf{A}\leftrightarrow\mathbf{B}$ and say that $\mathbf{A}$ and $\mathbf{B}$ are homomorphically equivalent.
For structures $A$ and $B$, we let $\hom(A,B)$ denote the number of homomorphisms from $A$ to $B$. We also use $\hom_{\mathbb{N}}(A,B)$ to denote this same number. We then define
$\hom_{\mathbb{B}}(A,B)\coloneqq\begin{cases}0&\hskip 6.0pt\text{if}\hskip 6.0pt\hom(A,B)=0\\ 1&\hskip 6.0pt\text{if}\hskip 6.0pt\hom(A,B)>0\end{cases}$
where $\mathbb{B}$ denotes the Boolean semiring. For a class $\mathcal{C}$ of structures, the left homomorphism profile of $A$ restricted to $\mathcal{C}$ is the tuple $(\hom(C,A))_{C\in\mathcal{C}}$. Similarly, we can define the right homomorphism profile and the homomorphism profile over $\mathbb{B}$. When we speak of a class of structures, we will always assume it is closed under isomorphism.
Properties of structures.
The incidence multigraph $\operatorname{inc}(\mathbf{A})$ of a structure $\mathbf{A}=(A,(R^{\mathbf{A}})_{R\in\tau})$ is the bipartite multigraph whose parts are the sets $A$ and $\operatorname{facts}(\mathbf{A})=\{(R,\mathbf{t}):R\in\tau\text{ and }\mathbf{t}\in R^{\mathbf{A}}\}$, and there is an edge between $a$ and $(R,\mathbf{t})$ for each entry in $\mathbf{t}$ that is $a$. We say that $\mathbf{A}$ is connected if $\operatorname{inc}(\mathbf{A})$ is connected, and we say $\mathbf{A}$ is acyclic if $\operatorname{inc}\mathbf{A}$ is acyclic. In particular, if an element appears multiple times in a fact, then the structure is not acyclic. This notion of acyclicity is also known as Berge-acyclicity. The number $c(\mathbf{A})$ of components of $\mathbf{A}$ is the maximal $n$ such that $A$ can be written as the disjoint union of $n$ structures.
For given structures $A$ and $B$ we let $A\oplus B$ denote their disjoint union and for a natural number $m$ and a structure $H$ we write
$m\cdot H\coloneqq\underbrace{H\oplus\ldots\oplus H}_{m-\text{times}}.$
Digraphs, walks, and homomorphisms to cycles.
A digraph (or a directed graph) is a structure with exactly one binary relation. So it is a pair $(V,R)$ where $R\subseteq V\times V$. A simple graph is an undirected graph without loops. We let $\mathsf{C}_{n}$ denote the directed cycle of length $n$:
$\mathsf{C}_{n}\coloneqq(\{0,\ldots,n-1\},\{(0,1),\ldots,(n-2,n-1),(n-1,0)\})$
and $\mathsf{P}_{n}$ denote the directed path of length $n$:
$\mathsf{P}_{n}\coloneqq(\{0,\ldots,n\},\{(0,1),\ldots,(n-1,n)\})$
For our discussion of adaptive left query algorithms, walks in directed graphs play a large role. We will thus review the definition of an oriented walk and some related concepts.
Definition 2.1.
Let $A=(V,R)$ be a digraph.
  1. (i)
    An oriented walk in $A$ is a sequence $(a_{0},r_{0},a_{1},r_{1},\ldots,a_{n},r_{n},a_{n+1})$ where $a_{i}\in V$ for each $i\in\{0,\ldots,n\}$, $r_{i}\in\{-,+\}$ and if $r_{i}=+$ then $(a_{i},a_{i+1})\in R$ while if $r_{i}=-$ then $(a_{i+1},a_{i})\in R$.
  2. (ii)
    The net length of an oriented walk $W=(a_{0},r_{0},a_{1},r_{1},\ldots,a_{n},r_{n},a_{n+1})$ is defined as:
[…]
  3. (iii)
    The oriented walk $(a_{0},r_{0},a_{1},r_{1},\ldots,a_{n},r_{n},a_{n+1})$ is said to be closed if $a_{0}=a_{n+1}$.
  4. (iv)
    A closed oriented walk $(a_{0},r_{0},a_{1},r_{1},\ldots,a_{n},r_{n},a_{0})$ is called an oriented cycle if $a_{i}\neq a_{j}$ for all $i\neq j$ and $n\geq 0$.
  5. (v)
    A directed walk is an oriented walk $(a_{0},r_{0},\ldots,a_{n})$ such that $r_{i}=+$ for each $i$. Closed directed walks and directed cycles are defined analogously.
An important thing to note is that oriented walks and their net lengths are preserved under homomorphisms. So if $(a_{0},r_{0},a_{1},r_{1},\ldots,a_{n},r_{n},a_{n+1})$ is an oriented walk in $A$ and $\varphi:A\to B$ is a homomorphism then
$(\varphi(a_{0}),r_{0},\varphi(a_{1}),r_{1},\ldots,\varphi(a_{n}),r_{n},\varphi(a_{n+1}))$
is an oriented walk in $B$ with the same net length.
For our proof of the existence of a class that is not decided by an adaptive left unbounded query algorithm, we take advantage of the fact that the existence of a homomorphism to a directed cycle can be described in an easy way.
Lemma 2.2 (Corollary 1.17 in [hell2004graphs]).
Let $A$ be a digraph and $n$ be a positive natural number. We have $A\to\mathsf{C}_{n}$ if and only if $n$ divides the net length of every closed oriented walk in $A$.
We use $\gcd$ to denote the greatest common divisor of a collection of natural numbers. For every finite digraph, we can define a parameter very related to the condition in Lemma 2.2.
Definition 2.3.
For a finite digraph $A$ we define
$\gamma(A)\coloneqq\gcd(l_{1},\ldots,l_{k})$
where $l_{1},\ldots,l_{k}$ is a listing of the net lengths of all cycles of positive net length in $A$.
Note, here we use the convention $\gcd(\varnothing)=0$. For integers $n,m$ we write $n\mid m$ if $n$ divides $m$, and $n\nmid m$ otherwise. The next proposition relates this parameter to the condition in Lemma 2.2.
Proposition 2.4.
Let $A$ be a digraph and $n$ be a positive integer. We have $A\to\mathsf{C}_{n}$ if and only if $n\mid\gamma(A)$.
Conjunctive queries.
A Boolean conjunctive query (CQ) is a first-order logic (FO) sentence of the form $q=\exists\textbf{x}(\bigwedge_{i}\alpha_{i})$ such that each $\alpha_{i}$ is of the form $R(y_{1},\ldots,y_{n})$ where each $y_{i}$ is among the variables in $\mathbf{x}$. It is well known that every such CQ has an associated “canonical structure” $A_{q}$ such that for every structure $B$ we have $B\vDash q$ if and only if $A_{q}\to B$. Similarly, vice versa, for every (finite) structure $A$, there is a Boolean CQ $q_{A}$ such that for every structure $B$ we have $A\to B$ if and only if $B\vDash q_{A}$. A CQ is said to be Berge-acyclic if its canonical structure is acyclic. A (Boolean) union of conjunctive queries (UCQ) is a finite disjunction of (Boolean) CQs. Every Boolean UCQ is a positive existential FO sentence, and, conversely, every positive existential FO sentence is equivalent to a UCQ.
3 Definitions of Query Algorithms
First, we review the definition of non-adaptive query algorithms.
Definition 3.1.
Let $\mathcal{C}$ be a class of relational structures with the same signature, let $\mathcal{K}\in\{\mathbb{N},\mathbb{B}\}$, and let $k$ be a positive integer.
  • A non-adaptive left $k$-query algorithm over $\mathcal{K}$ for $\mathcal{C}$ is a pair $(\mathcal{F},X)$, where $\mathcal{F}=(F_{1},\ldots,F_{k})$ is a tuple of relational structures with the same signature as the structures in $\mathcal{C}$, and $X$ is a set of $k$-tuples over $\mathcal{K}$, such that for all structures $D$, we have $D\in\mathcal{C}$ if and only if
    $(\hom_{\mathcal{K}}(F_{1},D),\hom_{\mathcal{K}}(F_{2},D),\ldots,\hom_{\mathcal{K}}(F_{k},D))\in X$
  • A non-adaptive right $k$-query algorithm over $\mathcal{K}$ for $\mathcal{C}$ is a pair $(\mathcal{F},X)$, where $\mathcal{F}=(F_{1},\ldots,F_{k})$ is a tuple of relational structures with the same signature as the structures in $\mathcal{C}$, and $X$ is a set of $k$-tuples over $K$, such that for all structures $D$, we have that $D\in\mathcal{C}$ if and only if
    $(\hom_{\mathcal{K}}(D,F_{1}),\hom_{\mathcal{K}}(D,F_{2}),\ldots,\hom_{\mathcal{K}}(D,F_{k}))\in X.$
Example 3.2.
Let $\mathcal{F}\coloneqq\{F\}$ where $F=(\{0\},\varnothing)$ is the singleton digraph with no edges, and let $X=\{3\}$. Then the non-adaptive left 1-query algorithm $(\mathcal{F},X)$ over $\mathbb{N}$ decides the class $\mathcal{C}$ of digraphs that have exactly 3 vertices. It is not difficult to see that $\mathcal{C}$ does not admit a non-adaptive left query algorithm over $\mathbb{B}$. Indeed, this follows from the fact that the class is not closed under homomorphic equivalence. A query algorithm over $\mathbb{B}$ cannot distinguish between structures that are homomorphically equivalent, since if $A\leftrightarrow B$ then $D\to A$ if and only if $D\to B$ for any $D$.
The example above shows that query algorithms over $\mathbb{N}$ have more expressive power than query algorithms over $\mathbb{B}$ in general. However, for classes closed under homomorphic equivalence, the same does not hold, as shown by ten Cate et al. in the following theorem.
Theorem 3.3 (Corollary 5.6 in [ten2024homomorphism]).
Let $\mathcal{C}$ be a class of structures that is closed under homomorphic equivalence. Then the following are equivalent:
  • $\mathcal{C}$ admits a non-adaptive left $k$-query algorithm over $\mathbb{N}$ for some $k$.
  • $\mathcal{C}$ admits a non-adaptive left $k$-query algorithm over $\mathbb{B}$ for some $k$.
  • $\mathcal{C}$ is FO-definable.
We now formulate the definition of an adaptive unbounded query algorithm. Adaptive query algorithms were defined in Chen et al. [chen_et_al:LIPIcs.MFCS.2022.32]. The definition there is limited in the sense that it only considers query algorithms whose number of queries is bounded by some constant.
For a set $\Sigma$, we let $\Sigma^{<\omega}$ denote the set of finite strings with alphabet $\Sigma$.
  • For $\sigma,\sigma^{\prime}\in\Sigma^{<\omega}$, we write $\sigma^{\prime}\sqsubseteq\sigma$ if $\sigma^{\prime}$ is an initial segment of $\sigma$.
  • We let $\sigma\bullet\sigma^{\prime}$ denote the concatenation of the strings $\sigma,\sigma^{\prime}$ (we also use this notation if $\sigma^{\prime}$ is an element of $\Sigma$).
  • If $\sigma_{0}\sqsubseteq\sigma_{1}\sqsubseteq\sigma_{2}\sqsubseteq\ldots$ is a sequence, then we let $\bigsqcup_{n\geq 0}\sigma_{n}$ denote the (possibly infinite) string $\sigma$ with length $m\coloneqq\sup_{n\geq 0}\absolutevalue{\sigma_{n}}$ | such that for $i\leq m$, the $i$-th letter of $\sigma$ is the $i$-th letter of $\sigma_{n}$ for any $n$ such that |$\absolutevalue{\sigma_{n}}\geq i$.
A set $\mathcal{T}\subseteq\Sigma^{<\omega}$ that is closed under initial segments (so if $\sigma\in\mathcal{T}$ and $\sigma^{\prime}\sqsubseteq\sigma$ then $\sigma^{\prime}\in\mathcal{T}$) is called a subtree of $\Sigma^{<\omega}$. An element $\sigma$ of such a subtree $\mathcal{T}$ is called a leaf if for every $\sigma^{\prime}\in\mathcal{T}$ such that $\sigma\sqsubseteq\sigma^{\prime}$ we have $\sigma^{\prime}=\sigma$.
We are now ready to define adaptive left unbounded query algorithms.
Definition 3.4.
Let $\mathcal{K}\in\{\mathbb{N},\mathbb{B}\}$.
  1. (i)
    An adaptive left unbounded query algorithm over $\mathcal{K}$ for structures with signature $\tau$ is a function
    $G:\mathcal{T}\to\mathsf{FIN}(\tau)\cup\{\textsf{YES},\textsf{NO}\}$
    where $\mathcal{T}$ is a subtree of $\mathcal{K}^{<\omega}$ and $G(\sigma)\in\{\textsf{YES},\textsf{NO}\}$ if and only if $\sigma$ is a leaf in $\mathcal{T}$.
  2. (ii)
    For $A\in\mathsf{FIN}(\tau)$, the computation path of an adaptive left unbounded query algorithm $G$ on $A$ is the string defined by $\sigma(A,G)\coloneqq\bigsqcup_{n\geq 0}\sigma_{n}$ where
    $\sigma_{0}$ $=\varepsilon$
    $\sigma_{n+1}$ =$=\begin{cases}\sigma_{n}&\hskip 6.0pt\text{if}\hskip 6.0ptG(\sigma_{n})\in\{\textsf{YES},\textsf{NO}\}\\ \sigma_{n}\bullet\hom_{\mathcal{K}}(G(\sigma_{n}),A)&\hskip 6.0pt\text{if}\hskip 6.0ptG(\sigma_{n})\in\mathsf{FIN}(\tau)\end{cases}.$
    The algorithm $G$ halts on input $A$ if $\sigma(A,G)$ is finite.
  3. (iii)
    If $G$ halts on all $A\in\mathsf{FIN}(\tau)$, then $G$ is said to be total.
  4. (iv)
    If $G$ is total, we say that it decides the class
    $\mathcal{C}\coloneqq\{A\in\mathsf{FIN}(\tau):G(\sigma(A,G))=\mathsf{YES}\}.$
  5. (v)
    We say that $G$ is an adaptive left $k$-query algorithm if for every $A$, $\sigma(A,G)$ has length at most $k$. We say $G$ is bounded if it is an adaptive left $k$-query algorithm for some $k$.
The notion of an adaptive right unbounded/bounded query algorithm $G$ is defined analogously by appending $\hom_{\mathcal{K}}(A,G(\sigma_{n}))$ instead of $\hom_{\mathcal{K}}(G(\sigma_{n}),A)$ in the computation path.
We will only consider total adaptive query algorithms, so when we speak of adaptive query algorithms we always assume it is total.
Example 3.5.
Let $G$ be an adaptive left query algorithm over $\mathbb{N}$ for digraphs defined with the following:
$G(\sigma)\coloneqq\begin{cases}(\{0\},\varnothing)&\hskip 6.0pt\text{if}\hskip 6.0pt\sigma=\varepsilon\\ \mathsf{P}_{n}&\hskip 6.0pt\text{if}\hskip 6.0pt\sigma=`n\textrm{'}\\ \mathsf{YES}&\hskip 6.0pt\text{if}\hskip 6.0pt\sigma=`a_{0}a_{1}\textrm{'}\hskip 6.0pt\text{and}\hskip 6.0pta_{1}\neq 0\\ \mathsf{NO}&\hskip 6.0pt\text{if}\hskip 6.0pt\sigma=`a_{0}a_{1}\textrm{'}\hskip 6.0pt\text{and}\hskip 6.0pta_{1}=0\end{cases}.$
The corresponding tree $\mathcal{T}\subseteq\mathbb{N}^{<\omega}$ would here be the set of all strings of length at most 2.
We can describe a run of the algorithm on a target structure $A$ as follows: First, the algorithm queries for the number of homomorphisms from the singleton digraph with no edges to $A$. Let $n$ denote the output of that query. It then queries for $\hom_{\mathbb{N}}(\mathsf{P}_{n},A)$. If the result is positive, it accepts; otherwise, it rejects.
To see what class this algorithm decides, we first note that the number of homomorphisms from the singleton digraph with no edges to $A$ is precisely the number of vertices of $A$. We also have that there exists a homomorphism from $\mathsf{P}_{n}$ to $A$ if and only if $A$ has a directed walk of length $n$. Since $n$ is the size of $A$, we see that such a walk has to visit the same vertex twice. We thus get that such a walk exists if and only if $A$ has a directed cycle. The algorithm, therefore, decides the class of digraphs that contain a directed cycle.
Example 3.6.
In this example, we show that every class $\mathcal{C}$ of structures is decided by an adaptive left unbounded query algorithm over $\mathbb{N}$.
The algorithm first queries for the size of the structure. Then it queries the number of homomorphisms from every structure that is not larger than it. By the proof of Lovász’s theorem (see, for example, Section III of [atserias2021expressive] for a proof of this), the algorithm has now distinguished the structure (up to isomorphism) and can thus classify it correctly.
To formally define the algorithm, let $\tau$ be the signature of the structures in $\mathcal{C}$. We fix an enumeration $A_{1},A_{2},A_{3},\ldots$ of all structures of signature $\tau$ such that for each $n\in\mathbb{N}$ there exists a number $r_{n}$ such that $A_{1},\ldots,A_{r_{n}}$ is an enumeration of all of the structures of size at most $n$. Then the algorithm can be defined with:
$G(\sigma)\coloneqq\begin{cases}(\{0\},\varnothing)&\hskip 6.0pt\text{if}\hskip 6.0pt\sigma=\varepsilon\\ A_{k+1}&\hskip 6.0pt\text{if}\hskip 6.0pt\sigma=`na_{1}a_{2}\ldots a_{k}\textrm{'}\text{ and }k<r_{n}\\ \textsf{YES}&\hskip 6.0pt\text{if}\hskip 6.0pt\sigma=`na_{1}a_{2}\ldots a_{r_{n}}\textrm{'}\text{ and the unique $A$ s.t. $\hom_{\mathbb{N}}(A_{i},A)=a_{i}$}\\ &\hskip 17.0pt\text{for each $1\leq i\leq r_{n}$ satisfies $A\in\mathcal{C}$}\\ \textsf{NO}&\hskip 6.0pt\text{if}\hskip 6.0pt\sigma=`na_{1}a_{2}\ldots a_{r_{n}}\textrm{'}\text{ and the unique $A$ s.t. $\hom_{\mathbb{N}}(A_{i},A)=a_{i}$}\\ &\hskip 17.0pt\text{for each $1\leq i\leq r_{n}$ satisfies $A\notin\mathcal{C}$}\end{cases}$
4 Query Algorithms over $\mathbb{N}$
4.1 Unboundedness helps for adaptive left query algorithms
A starting point for our work in this section is the following result by Chen et al:
Theorem 4.1 (Theorem 8.2 in [chen_et_al:LIPIcs.MFCS.2022.32]).
Every class of simple graphs can be decided by an adaptive left 3-query algorithm over $\mathbb{N}$.
This result invites the question of what happens in the general case, for arbitrary relational structures. We prove that the same does not hold in this broader setting. We even show that there is a class of directed graphs that is not decided by an adaptive $k$-query algorithm over $\mathbb{N}$, for any $k$. To do this, we utilize the symmetry of directed cycles to prove that they are hard to distinguish from each other using left homomorphism queries.
We begin with a lemma that shows that the homomorphism count from a structure to a disjoint union of cycles can be described simply.
Lemma 4.2.
For every digraph $A$ and all positive integers $n,m$ we have
$\hom(A,m\cdot\mathsf{C}_{n})\coloneqq\begin{cases}0&\hskip 6.0pt\text{if}\hskip 6.0ptn\nmid\gamma(A)\\ (m\cdot n)^{c(A)}&\hskip 6.0pt\text{if}\hskip 6.0ptn\mid\gamma(A)\end{cases}.$
The proof of this theorem is in Appendix LABEL:appendix:basic-digraphs. The proof of this theorem can be intuitively explained as follows: In disjoint unions of directed cycles, each vertex has in-degree 1 and out-degree 1. For a homomorphism from $A\to m\cdot\mathsf{C}_{n}$, the whole map is therefore decided by the value of the map in one vertex from each component of $A$. Since there are $m\cdot n$ ways to pick the value of a homomorphism on that initial vertex (if such a homomorphism exists), the lemma follows.
We say that two classes $\mathcal{A},\mathcal{B}$ of structures are separated by a query algorithm if it accepts all structures in $\mathcal{A}$ and rejects all structures in $\mathcal{B}$, or vice versa.
The importance of the above lemma is in the fact that it shows that for a given structure $m\cdot\mathsf{C}_{n}$ there are only two possible outcomes for $\hom(F,m\cdot\mathsf{C}_{n})$. This limits what can be done using few queries, as is shown in the following theorem:
Theorem 4.3.
The class $\mathcal{D}_{n}\coloneqq\Big{\{}2^{n-m}\cdot\mathsf{C}_{2^{m}}:0\leq m\leq n\hskip 6.0pt\text{and}\hskip 6.0ptm\text{ is even }\Big{\}}$ can be separated from the class $\mathcal{D}_{n}^{\prime}\coloneqq\Big{\{}2^{n-m}\cdot\mathsf{C}_{2^{m}}:0\leq m\leq n\hskip 6.0pt\text{and}\hskip 6.0ptm\text{ is odd }\Big{\}}$ by a non-adaptive left $k$-query algorithm over $\mathbb{N}$ if and only if $k\geq n$.
Proof 4.4.
By Lemma 4.2 we have that for every digraph $F$ and every $2^{n-m}\cdot\mathsf{C}_{2^{m}}\in\mathcal{D}_{n}\cup\mathcal{D}_{n}^{\prime}$:
$\hom(F,2^{n-m}\cdot\mathsf{C}_{2^{m}})$ =$=\begin{cases}0&\hskip 6.0pt\text{if}\hskip 6.0pt2^{m}\nmid\gamma(F)\\ (2^{n})^{c(F)}&\hskip 6.0pt\text{if}\hskip 6.0pt2^{m}\mid\gamma(F)\end{cases}$
=$=\begin{cases}0&\hskip 6.0pt\text{if}\hskip 6.0pt\nu_{2}(\gamma(F))<m\\ (2^{n})^{c(F)}&\hskip 6.0pt\text{if}\hskip 6.0pt\nu_{2}(\gamma(F))\geq m\end{cases}$
where $\nu_{2}(k)$ denotes the largest positive integer $r$ such that $2^{r}\mid k$ if $k\neq 0$, if $k=0$ then $\nu_{2}(k)=+\infty$. It follows from this that to have $\hom(F,2^{n-m}\cdot\mathsf{C}_{2^{m}})\neq\hom(F,2^{n-(m+1)}\cdot\mathsf{C}_{2^{m+1}})$ we must have $\nu_{2}(\gamma(F))=m+1$. To distinguish between $2^{n-m}\cdot\mathsf{C}_{2^{m}}$ and $2^{n-(m+1)}\cdot\mathsf{C}_{2^{m+1}}$ for every $m\in\{0,\ldots,n-1\}$ the algorithm then needs to query some $F_{m}$ with $\nu_{2}(\gamma(F_{m}))=m$ for each $m\in\{0,\ldots,n-1\}$. The algorithm, therefore, needs at least $n$ queries to separate the classes. It is clear that $n$ non-adaptive queries suffice: the algorithm can query $\mathsf{C}_{2^{m}}$ for each $m\in\{0,\ldots,n-1\}$. Every structure in $\mathcal{D}_{n}\cup\mathcal{D}_{n}^{\prime}$ yields a unique outcome from the collection of these queries, and thus they can be classified correctly.
It is worth noting that homomorphism existence queries suffice here, and hence $\mathcal{D}_{n}$ and $\mathcal{D}^{\prime}_{n}$ are already separated by a non-adaptive left $n$-query algorithm over $\mathbb{B}$. ∎
Corollary 4.5.
The class $\mathcal{D}_{n}$ can be separated from the class $\mathcal{D}_{n}^{\prime}$ by an adaptive left $k$-query algorithm over $\mathbb{N}$ if and only if $k\geq\log(n+1)$.